OpenStudy (anonymous):

Erik needs a plumber and must decide between two companies. For a service visit, Plumber A charges $85 to send a plumber, plus $20/h. Plumber B charges $60 to send a plumber, plus $30/h. The system that models this situation is given, where c is the cost of hiring a plumber and h is the number of hours the plumber works. The solution to the system is (2.5, 135). What is the correct interpretation of the solution of this system? (this is the first part )

OpenStudy (anonymous):

c=85+20h c=60+30h A.Plumber A will charge more when he works 2.5 h, charging $135. B.Plumber B will charge more when she works 2.5 h, charging $135. C.Plumber A and Plumber B will both charge $135 if they each work 2.5 h. D.Plumber A and Plumber B will both charge $2.50 if they each work 135 h. please help will fan and medal